A gas undergoes an isothermal expansion form V 1-3.7 to lowed by isobanc p cs compression. Ifp1-2.8 m p2 1.50 m → ? 2, calculate the work done by gas during the expansion.
Answer #1

Since for isothermal process,

PV=nRT for n moles of gas

P_1V_1=nRT

T is same throughout isothermal process

P_1=2.8 atm, V_1=3.7LP_1=2.8 atm=2.8*101325 Pa= 283710 Pa

V_1=3.7L=3.7*10^{-3}m^3

so nRT=P_1V_1=283710*3.7*10^{-3}=1049.727J

so work done during isothermal expansion

W=nRT ln\frac{V_2}{V_1}=nRT ln\frac{P_1}{P_2}

W=1049.727 ln\frac{2.8}{1.5}=1049.727 *2.303 log_{10} \left (\frac{2.8}{1.5} \right )

W=1049.727*2.303*0.27=652.73J